# AutoNeural: Co-Designing Vision–Language Models for NPU Inference

### v1

@@ 1 Introduction
Vision–language models (VLMs) [1, 2, 3, 4, 5] have fundamentally advanced the capability to jointly reason over images and text, achieving remarkable success in recognition, grounding, captioning, and visual instruction following. The dominant paradigm typically grafts a pretrained Vision Transformer (ViT) onto a decoder-only Large Language Model (LLM) via a projection layer, aligning the modalities through visual instruction tuning. While this GPU-centric design yields high accuracy at large resolutions, it introduces systemic inefficiencies when deployed on edge devices. Specifically, it suffers from: (i) high inference latency, dominated by the heavy computational cost of vision encoding and the autoregressive generation of the LLM (affecting both time-to-first-token and time-per-token); and (ii) inherent brittleness under low-precision execution, owing to the attention-heavy architectures in both vision and language components that are sensitive to quantization.
@@
Neural Processing Units (NPUs) [6, 7, 8] have emerged as the default compute substrate for mobile and edge intelligence, delivering high TOPS and energy efficiency through specialized integer operators and on-chip SRAM. However, state-of-the-art VLMs optimized for floating-point GPU execution often degrade significantly under NPU-friendly INT- quantization constraints. We identify two root causes for this hardware–software mismatch. First, Vision Transformers (ViTs) [9, 10] exhibit quantization brittleness—activations and weights in multi-head attention and RMSNorm [11] paths often possess outlier distributions that are sensitive to low-precision scaling, leading to sharp accuracy collapse at INT8/16. Second, Transformer-based language backbones suffer from memory I/O bottlenecks. The autoregressive generation process necessitates repeated Key-Value (KV) cache access, creating memory-bound operations that saturate on-chip bandwidth. Consequently, the NPU’s compute units stall despite high nominal TOPS, inflating latency and reducing end-to-end throughput.
@@
We introduce AutoNeural, an NPU-native VLM architecture that fundamentally rethinks vision encoding and language modeling for edge efficiency. For the vision encoder, we depart from standard ViTs in favor of a MobileNet-style architecture utilizing depthwise separable convolutions. This design eliminates the quadratic complexity and quantization brittleness of global attention, providing strong inductive biases for feature extraction while maintaining bounded activation distributions inherently stable for INT8/16 inference. For the language backbone, we adopt a hybrid Transformer-SSM architecture (inspired by Liquid AI principles), which interleaves Transformer layers with efficient gated convolutions based on structural state-space models. This topology offers linear-time complexity and compact state representations, obviating the need for explicit KV caching and reducing memory I/O by up to 60% during generation. To our knowledge, this is the first framework to unify convolution-based visual stability with linear-complexity state-space modeling, effectively decoupling multi-modal reasoning capability from the heavy computational burden of global attention. We couple this architecture with an NPU-aware training recipe that integrates quantization-aware fine-tuning (QAT) [12, 13, 14], mixed-precision constraints, and hardware-aligned calibration.
@@ 2 Related Work
The automotive industry has increasingly adopted vision–language models to enable advanced driver assistance systems (ADAS), in-car assistants, and autonomous driving capabilities [15, 16]. Recent work has explored the application of large multimodal models in automotive contexts, ranging from scenario generation to real-time perception. For instance, multimodal LLMs [17, 1]have been applied to autonomous driving tasks such as scenario generation and perception, demonstrating the potential of integrating visual and textual reasoning for improved situational awareness[18, 19, 20, 21, 22]. However, most of these systems rely on GPU-based inference and have not been optimized for the power and latency constraints of in-vehicle NPUs. The deployment of such models in production vehicles requires careful co-design with edge hardware, a gap our work directly addresses. While existing automotive multimodal systems often compromise on model capacity or resolution to meet latency budgets, our MobileNet-Hybrid Transformer-SSM architecture demonstrates that hardware-aware design enables both high accuracy and low latency.
@@
Neural Processing Units from vendors such as Qualcomm [23, 24], MediaTek [25], and Apple have become the primary compute substrate for on-device AI across mobile and automotive platforms. Recent efforts have focused on adapting large multimodal models for NPU deployment through quantization [26, 27], pruning [28], and operator fusion [29]. For example, work on Qualcomm platforms [30, 31, 32, 33] has demonstrated that quantization-aware training and low-rank factorization can accelerate inference while maintaining acceptable accuracy. Similarly, frameworks like MindVL [34] have explored efficient training pipelines on Ascend NPUs, integrating distributed data loading and system-level scheduling. MiniCPM-V [35, 36] represents another line of work targeting edge deployment through aggressive model compression, achieving strong performance on mobile devices. However, these approaches primarily focus on post-hoc optimization of GPU-first architectures rather than fundamental architectural redesign. In contrast, our work rethinks both the vision encoder and language backbone to match NPU operator sets and memory hierarchies, yielding more substantial latency and efficiency gains.
@@
The choice of vision encoder critically impacts both accuracy and deployment efficiency in multimodal systems. While Vision Transformers have dominated recent VLM architectures due to their strong performance, their global attention mechanisms and quantization brittleness pose significant challenges for on-device deployment. Recent work has explored alternative encoder designs optimized for edge scenarios. MobileNet architectures [37, 38, 39, 40], originally designed for efficient image classification, employ depthwise separable convolutions that reduce parameter count and computation while maintaining strong inductive biases for visual features. Notably, PaliGemma [41, 42], a recent 3B parameter VLM from Google, demonstrates that carefully designed smaller vision encoders can achieve competitive transfer performance. Beyond vision encoders, recent advances in sequence modeling have introduced State Space Models (SSMs) [43, 44, 45, 46]as efficient alternatives to Transformers for language modeling. SSMs achieve linear-time complexity through selective state spaces and eliminate the need for explicit key-value caching, making them particularly attractive for memory-constrained NPU deployment. Building on these principles, Liquid Foundation Models [47, 48, 49] demonstrate efficient sequence modeling architectures that combine linear-time complexity with strong representational capacity through liquid structural state-space modeling. Our work adopts the Liquid AI architecture as our language backbone and systematically pairs it with a MobileNet-based encoder optimized for NPU-aware quantization. By replacing global attention in both the vision and language components, we address the dual bottlenecks of quantization brittleness and memory I/O that plague existing VLM architectures on NPUs.
@@ 3.1 Model Architecture
Our vision encoder builds on MobileNetV5, exploiting depthwise separable convolutions for efficient feature extraction with bounded activations, and is initialized from the Gemma 3n-E4B vision checkpoint to align with the multimodal design of Gemma 3n [50, 51]. It consumes inputs and produces a spatial feature map that is flattened into visual tokens. In contrast to ViTs with global self-attention, we employ a hierarchical stack of inverted residual (IR) blocks that naturally yields multi-scale features while preserving local receptive fields; this local-first design reduces quantization sensitivity, as convolution + normalization maintain stabler activation ranges than LayerNorm and multi-head attention, improving INT8/16 robustness. The network begins with a stride-2 stem, followed by four stages in which stride-2 IR blocks downsample and stride-1 IR blocks refine. To inject long-range context at bounded cost, the late, low-resolution stages interleave sparse multi-query attention (MQA) bottlenecks. We aggregate information via a Multi-Scale Fusion Adapter (MSFA): the last two stage outputs are upsampled to the finest tapped resolution, channel-concatenated, and processed by a universal inverted residual layer (pointwisedepthwisepointwise) with GELU (tanh) activations and RMSNorm; a stride-3 average pool yields a fused tensor. This spatial feature map is then flattened into visual tokens, where each of the spatial locations becomes a token with -dimensional features, balancing fidelity and LLM prefill cost while preserving accuracy and latency benefits.
@@ 3.4 Cockpit AI and automotive applications
Modern automotive cockpits are evolving from traditional human-machine interfaces into intelligent multimodal assistants that understand both visual and textual contexts [52, 53]. Edge-deployed multimodal models enable diverse interaction scenarios critical for driving safety and user experience. These include real-time interpretation of dashboard indicators and warning lights with actionable recommendations based on vehicle sensor data. Scene understanding for navigation assistance under challenging conditions, such as nighttime or adverse weather when visibility is compromised, and interpretation of traffic signs and road conditions through external cameras help prevent navigation errors. The system also supports analysis of documents captured via in-cabin cameras, multimodal interaction combining gesture recognition with voice commands for hands-free operation, and driver state monitoring to detect fatigue or anxiety and provide timely interventions through environmental adjustments.
@@ 4.1 Training setup
We utilize the Infinity-MM dataset [54], a large-scale, high-quality multimodal instruction dataset designed for comprehensive vision–language training. Infinity-MM contains approximately 44.8 million instruction-following samples spanning diverse visual understanding capabilities. The dataset is carefully curated to balance multiple task categories including general VQA covering everyday visual reasoning, object recognition, and scene understanding; document understanding with forms, receipts, scientific papers, and business documents; chart and diagram reasoning; OCR-centric tasks; multi-turn conversations; and specialized domains such as medical imaging, autonomous driving scenarios, and industrial inspection. Each sample consists of an image, an instruction sequence, and reference responses. The dataset emphasizes high-quality, detailed responses that require multi-step reasoning rather than simple label recognition, making it particularly suitable for training models that need to provide thorough explanations—critical for automotive safety applications where users must understand the model’s reasoning process.
@@ 4.2 Comparison with other models
We evaluate our NPU-native architecture against state-of-the-art vision–language models across five diverse multimodal benchmarks. Our comparison includes recent VLMs spanning different architectural families and parameter scales: InternVL2 (1B and 2B variants) representing ViT-Transformer architectures optimized for GPU deployment, and Qwen2-VL (2B and 3.75B variants) [55] demonstrating strong general-purpose multimodal capabilities. We report performance of our AutoNeural model trained to Stage 3 of the Infinity-MM curriculum, which achieves the best balance between accuracy and training efficiency.
@@ Results.
| Model | Params | MMStar | Hallusion | MathVista | AI2D | OCR | Avg |
| Bench | MINI | TEST | Bench | ||||

### v2

@@ 1 Introduction
Vision–language models (VLMs) [chen2024omnivlmtokencompressedsubbillionparametervisionlanguage, marafioti2025smolvlm, zhu2025internvl3, zhang2024vision, ghosh2024exploring] have fundamentally advanced the capability to jointly reason over images and text, achieving remarkable success in recognition, grounding, captioning, and visual instruction following. The dominant paradigm typically grafts a pretrained Vision Transformer (ViT) onto a decoder-only Large Language Model (LLM) via a projection layer, aligning the modalities through visual instruction tuning. While this GPU-centric design yields high accuracy at large resolutions, it introduces systemic inefficiencies when deployed on edge devices. Specifically, it suffers from: (i) high inference latency, dominated by the heavy computational cost of vision encoding and the autoregressive generation of the LLM (affecting both time-to-first-token and time-per-token); and (ii) inherent brittleness under low-precision execution, owing to the attention-heavy architectures in both vision and language components that are sensitive to quantization.
@@
Neural Processing Units (NPUs) [tan2021efficient, 10472723, lee2021architecture] have emerged as the default compute substrate for mobile and edge intelligence, delivering high TOPS and energy efficiency through specialized integer operators and on-chip SRAM. However, state-of-the-art VLMs optimized for floating-point GPU execution often degrade significantly under NPU-friendly INT- quantization constraints. We identify two root causes for this hardware–software mismatch. First, Vision Transformers (ViTs) [khan2022transformers, chen2024comprehensive] exhibit quantization brittleness—activations and weights in multi-head attention and RMSNorm [xu2019understanding] paths often possess outlier distributions that are sensitive to low-precision scaling, leading to sharp accuracy collapse at INT8/16. Second, Transformer-based language backbones suffer from memory I/O bottlenecks. The autoregressive generation process necessitates repeated Key-Value (KV) cache access, creating memory-bound operations that saturate on-chip bandwidth. Consequently, the NPU’s compute units stall despite high nominal TOPS, inflating latency and reducing end-to-end throughput.
@@
We introduce AutoNeural, an NPU-native VLM architecture that fundamentally rethinks vision encoding and language modeling for edge efficiency. For the vision encoder, we depart from standard ViTs in favor of a MobileNet-style architecture utilizing depthwise separable convolutions. This design eliminates the quadratic complexity and quantization brittleness of global attention, providing strong inductive biases for feature extraction while maintaining bounded activation distributions inherently stable for INT8/16 inference. For the language backbone, we adopt a hybrid Transformer-SSM architecture (inspired by Liquid AI principles), which interleaves Transformer layers with efficient gated convolutions based on structural state-space models. This topology offers linear-time complexity and compact state representations, obviating the need for explicit KV caching and reducing memory I/O by up to 60% during generation. To our knowledge, this is the first framework to unify convolution-based visual stability with linear-complexity state-space modeling, effectively decoupling multi-modal reasoning capability from the heavy computational burden of global attention. We couple this architecture with an NPU-aware training recipe that integrates quantization-aware fine-tuning (QAT) [li2023loftq, bondarenko2024low, chen2025efficientqat], mixed-precision constraints, and hardware-aligned calibration.
@@ 2 Related Work
The automotive industry has increasingly adopted vision–language models to enable advanced driver assistance systems (ADAS), in-car assistants, and autonomous driving capabilities [guo2024vlm, huang2024vlm]. Recent work has explored the application of large multimodal models in automotive contexts, ranging from scenario generation to real-time perception. For instance, multimodal LLMs [chen2024octopusv3technicalreport, chen2024omnivlmtokencompressedsubbillionparametervisionlanguage]have been applied to autonomous driving tasks such as scenario generation and perception, demonstrating the potential of integrating visual and textual reasoning for improved situational awareness[chen-etal-2025-octopus, chen2024octopusv2ondevicelanguage, chen2024octopus, chen2024octoplannerondevicelanguagemodel, chen2024squidlongcontextnew]. However, most of these systems rely on GPU-based inference and have not been optimized for the power and latency constraints of in-vehicle NPUs. The deployment of such models in production vehicles requires careful co-design with edge hardware, a gap our work directly addresses. While existing automotive multimodal systems often compromise on model capacity or resolution to meet latency budgets, our MobileNet-Hybrid Transformer-SSM architecture demonstrates that hardware-aware design enables both high accuracy and low latency.
@@
Neural Processing Units from vendors such as Qualcomm [zhu2025edge, xu2025fast], MediaTek [park2023overviewing], and Apple have become the primary compute substrate for on-device AI across mobile and automotive platforms. Recent efforts have focused on adapting large multimodal models for NPU deployment through quantization [wei2024advances, tmamna2024pruning], pruning [dantas2024comprehensive], and operator fusion [salmani2025llm]. For example, work on Qualcomm platforms [federici2024efficient, rajaee2025locallookaheadguidanceverifierintheloop, hegde2025distillingmultimodallargelanguage, farhadzadeh2025loraxbridgingfoundationmodels] has demonstrated that quantization-aware training and low-rank factorization can accelerate inference while maintaining acceptable accuracy. Similarly, frameworks like MindVL [chen2025mindvl] have explored efficient training pipelines on Ascend NPUs, integrating distributed data loading and system-level scheduling. MiniCPM-V [yu2025minicpm, yao2024minicpm] represents another line of work targeting edge deployment through aggressive model compression, achieving strong performance on mobile devices. However, these approaches primarily focus on post-hoc optimization of GPU-first architectures rather than fundamental architectural redesign. In contrast, our work rethinks both the vision encoder and language backbone to match NPU operator sets and memory hierarchies, yielding more substantial latency and efficiency gains.
@@
The choice of vision encoder critically impacts both accuracy and deployment efficiency in multimodal systems. While Vision Transformers have dominated recent VLM architectures due to their strong performance, their global attention mechanisms and quantization brittleness pose significant challenges for on-device deployment. Recent work has explored alternative encoder designs optimized for edge scenarios. MobileNet architectures [chen2018enhanced, sinha2019thin, qin2018fd, khasoggi2019efficient], originally designed for efficient image classification, employ depthwise separable convolutions that reduce parameter count and computation while maintaining strong inductive biases for visual features. Notably, PaliGemma [beyer2024paligemma, steiner2024paligemma], a recent 3B parameter VLM from Google, demonstrates that carefully designed smaller vision encoders can achieve competitive transfer performance. Beyond vision encoders, recent advances in sequence modeling have introduced State Space Models (SSMs) [gu2021efficiently, gu2024mamba, dao2024transformersssmsgeneralizedmodels, gu2021combining]as efficient alternatives to Transformers for language modeling. SSMs achieve linear-time complexity through selective state spaces and eliminate the need for explicit key-value caching, making them particularly attractive for memory-constrained NPU deployment. Building on these principles, Liquid Foundation Models [wu2024liquid, smekal2024towards, poli2403mechanistic] demonstrate efficient sequence modeling architectures that combine linear-time complexity with strong representational capacity through liquid structural state-space modeling. Our work adopts the Liquid AI architecture as our language backbone and systematically pairs it with a MobileNet-based encoder optimized for NPU-aware quantization. By replacing global attention in both the vision and language components, we address the dual bottlenecks of quantization brittleness and memory I/O that plague existing VLM architectures on NPUs.
@@ 3.1 Model Architecture
Our vision encoder builds on MobileNetV5, exploiting depthwise separable convolutions for efficient feature extraction with bounded activations, and is initialized from the Gemma 3n-E4B vision checkpoint to align with the multimodal design of Gemma 3n [gemma_3n_2025, team2025gemma]. It consumes inputs and produces a spatial feature map that is flattened into visual tokens. In contrast to ViTs with global self-attention, we employ a hierarchical stack of inverted residual (IR) blocks that naturally yields multi-scale features while preserving local receptive fields; this local-first design reduces quantization sensitivity, as convolution + normalization maintain stabler activation ranges than LayerNorm and multi-head attention, improving INT8/16 robustness. The network begins with a stride-2 stem, followed by four stages in which stride-2 IR blocks downsample and stride-1 IR blocks refine. To inject long-range context at bounded cost, the late, low-resolution stages interleave sparse multi-query attention (MQA) bottlenecks. We aggregate information via a Multi-Scale Fusion Adapter (MSFA): the last two stage outputs are upsampled to the finest tapped resolution, channel-concatenated, and processed by a universal inverted residual layer (pointwisedepthwisepointwise) with GELU (tanh) activations and RMSNorm; a stride-3 average pool yields a fused tensor. This spatial feature map is then flattened into visual tokens, where each of the spatial locations becomes a token with -dimensional features, balancing fidelity and LLM prefill cost while preserving accuracy and latency benefits.
@@ 3.4 Cockpit AI and automotive applications
Modern automotive cockpits are evolving from traditional human-machine interfaces into intelligent multimodal assistants that understand both visual and textual contexts [zhang2021pilot, kovalyov2024intelligent]. Edge-deployed multimodal models enable diverse interaction scenarios critical for driving safety and user experience. These include real-time interpretation of dashboard indicators and warning lights with actionable recommendations based on vehicle sensor data. Scene understanding for navigation assistance under challenging conditions, such as nighttime or adverse weather when visibility is compromised, and interpretation of traffic signs and road conditions through external cameras help prevent navigation errors. The system also supports analysis of documents captured via in-cabin cameras, multimodal interaction combining gesture recognition with voice commands for hands-free operation, and driver state monitoring to detect fatigue or anxiety and provide timely interventions through environmental adjustments.
@@ 4.1 Training setup
We utilize the Infinity-MM dataset [gu2024infinitymmscalingmultimodalperformance], a large-scale, high-quality multimodal instruction dataset designed for comprehensive vision–language training. Infinity-MM contains approximately 44.8 million instruction-following samples spanning diverse visual understanding capabilities. The dataset is carefully curated to balance multiple task categories including general VQA covering everyday visual reasoning, object recognition, and scene understanding; document understanding with forms, receipts, scientific papers, and business documents; chart and diagram reasoning; OCR-centric tasks; multi-turn conversations; and specialized domains such as medical imaging, autonomous driving scenarios, and industrial inspection. Each sample consists of an image, an instruction sequence, and reference responses. The dataset emphasizes high-quality, detailed responses that require multi-step reasoning rather than simple label recognition, making it particularly suitable for training models that need to provide thorough explanations—critical for automotive safety applications where users must understand the model’s reasoning process.
@@ 4.2 Comparison with other models
We evaluate our NPU-native architecture against state-of-the-art vision–language models across five diverse multimodal benchmarks. Our comparison includes recent VLMs spanning different architectural families and parameter scales: InternVL2 (1B and 2B variants) representing ViT-Transformer architectures optimized for GPU deployment, and Qwen2-VL (2B and 3.75B variants) [wang2024qwen2] demonstrating strong general-purpose multimodal capabilities. We report performance of our AutoNeural model trained to Stage 3 of the Infinity-MM curriculum, which achieves the best balance between accuracy and training efficiency.
@@ Results.
| Model | Params | AI2D | Hallusion | MathVista | MMStar | OCR | Avg |
| TEST | Bench | MINI | Bench | ||||